Given that sec θ = 5/4 and tan θ < 0 , find the value of sin(2θ).
Answer by amarjeeth123(569)

sec θ = 5/4 and tanθ<0
This is possible iff theta lies in the fourth quadrant.
cos θ=4/5
sin θ=-3/5 because sin theta is negative in the fourth quadrant.
sin(2θ)=2sin θcos θ=2(-3/5)(4/5)=-24/25
Answer=-(24/25)